1. A steel tank at 20.97°C contains a gas under a pressure of 5.97 atm. For safety reasons, it is desired that the pressure not exceeded 8.5 atm. What is the maximum temperature (in °C) to which the gas in tank can be heated?



1
Expert's answer
2021-01-26T03:21:27-0500

"\frac{P_1}{T_1}" = "\frac{P_2}{T_2}"

"\frac{5.97atm}{20.97⁰C}" = "\frac{8.5atm}{T_2}"

T2 = 29.85⁰C